What is the NMR structure given the information provided? The NMR spectrum, IR spectrum, and molecular formula are the...

0-lnrV-z-E0WKaYVcUjA_bErsu4telJLhFkoNWRg

What is the NMR structure given the information provided?


The NMR spectrum, IR spectrum, and molecular formula are the only things that were provided for the question. I'm trying to figure out what the structure would be give the information.
Answer #1

Molecular formual CH15NO DBE =C+1 - H/2 +N/2 = 7+1 -15/2 + 1/2 = 8.5 - 7.5 = 1 IR shows that strong peak at 1640 cm-1 which i